OpenStudy (s117):

The solution to which inequality is shown? A. y + 5 ≤ 4 B. y + 5 < 4 C. y + 5 ≥ 4 D. y + 5 > 4

OpenStudy (anonymous):

\(<\) or \(>\) will have an open dot \(\le\) or \(\ge\) will have a closed dot

OpenStudy (anonymous):

So which one's can you cross out? @S117

OpenStudy (s117):

a and c so its between b and d

OpenStudy (anonymous):

Actually, we cross out B and D..because their signs will make an open dot. And there is a closed dot on the graph.

OpenStudy (s117):

oh yea

OpenStudy (s117):

@iGreen.

OpenStudy (anonymous):

Okay, which one do you think?

OpenStudy (anonymous):

\(>\) or \(\ge\) will point to the right. \(<\) or \(\le\) will point to the left.

OpenStudy (anonymous):

@S117 What do you think? A or C?

OpenStudy (s117):

c

OpenStudy (s117):

one more?

OpenStudy (anonymous):

Sure